# Contactless Fingerprint Recognition Using 3D Graph Matching

### v1

@@ Abstract
Contactless fingerprint is a newly developed type of fingerprint, and has gained lots of attention in recent fingerprint studies. However, most existing contactless fingerprint algorithms treat contactless fingerprints as 2D plain fingerprints, and utilize similar recognition methods as traditional contact-based 2D fingerprints. This recognition approach does not consider the modality difference between contactless and contact fingerprints, especially the intrinsic 3D characteristic of contactless fingerprints. This paper proposes a novel contactless fingerprint recognition algorithm that captures the revealed 3D feature of contactless fingerprints rather than the plain 2D feature. The proposed method first recovers 3D features from the input contactless fingerprint, including the 3D shape model and 3D fingerprint feature (minutiae, orientation, etc.). Then, a novel 3D graph matching is conducted in 3D space according to the extracted 3D feature. Our method captures the real 3D nature of contactless fingerprints as the whole feature extraction and matching algorithms are completed in real 3D space. Experiments results on contactless fingerprint databases show that the proposed method successfully improves the matching accuracy of contactless fingerprints. Exceptionally, our method performs stably across multiple poses of contactless fingerprints due to 3D graph matching, which is a great advantage compared to previous contactless fingerprint recognition algorithms.
Index Terms:
Contactless fingerprint, 3D fingerprint recognition, graph matching.
Figure 1: The pipeline of the proposed method. Our method first extracts 3D features from the input contactless 2D fingerprint, then uses the extracted 3D features to conduct graph matching directly in 3D space.
I Introduction
The fingerprint recognition system has become one of the most widely used biometric systems for person authentication and recognition tasks. To improve fingerprint recognition performances, researchers have developed lots of fingerprint algorithms in various aspects of fingerprints [1]. Among them, contactless fingerprint recognition has gained a lot of attention in recent studies [2].
@@ I Introduction
Compared with traditional contact-based fingerprint, which relies on pressing finger skin against sensors to gather fingerprint images, contactless fingerprint provides a touchless approach to acquiring fingerprint images through a specific imaging system [3]. Since the collection process is touch-free, the gathered fingerprints are of no skin distortion, which is a crucial problem in contact-based fingerprint matching [4]. The collection process is more hygienic and convenient, and can even be done with mobile phone cameras [5][6].
@@
Despite these advantages, contactless fingerprint recognition still meets challenges on modality differences with contact-based fingerprints [7]. The contactless fingerprint image is acquired by a camera-like imaging system [8][9] in 3D space, which differs from the traditional contact-based 2D sensors. The gathered contactless fingerprint is theoretically a projection from a live 3D finger to a 2D plane image. The finger surface curvature is incorporated into the imaging result, which leads to evident perspective distortion on contactless fingerprints [10][11]. The distortion causes changes in ridge orientation, period, and minutiae location, thus degrading recognition performances [12].
@@
To overcome perspective distortion and accurately match contactless fingerprints, researchers have proposed methods to deal with this issue. A significant approach is to treat perspective distortion similar to elastic skin distortion [13] in contact-based 2D fingerprints. A transformation network [14][15][16][17] is trained to fit the distortion, thus rectifying the fingerprint to the undistorted level. This approach is generally used in contactless-to-contact (CL2C) matching, since it treats a contactless fingerprint as a distorted 2D fingerprint, similar to a contact-based 2D fingerprint.
@@
As for contactless-to-contactless (CL2CL) matching, several studies have further considered the intrinsic 3D information of contactless fingerprints. As the contactless fingerprint is a projection of a 3D finger, a finger shape model [11][7][18] is recovered from the original contactless fingerprint. Then, all the contactless fingerprints are rotated to the same pose (usually the front pose) in 3D space to eliminate pose variations for improving the matching accuracy.
@@
Although significant progress has been made in contactless fingerprint recognition, the modality differences between contactless and contact fingerprints have yet to be fully studied. Most algorithms [14][15][19][16][17] treat contactless fingerprints as plain 2D fingerprint images, therefore neglecting the 3D information hidden in contactless fingerprints. Some studies further consider 3D information. However, the recovered 3D shape is only used for pose correction [10][11] or unwarping [20][18]. The rotated or unwarped contactless fingerprint is still treated as a 2D image for the following feature extraction and matching steps.
@@
In this paper, we propose a novel contactless fingerprint recognition that entirely uses 3D information in contactless fingerprints. As shown in Fig. 1, the proposed method takes a single contactless fingerprint as input, and then extracts its 3D feature, including the shape model and 3D minutiae feature. The extracted 3D features are finally used for 3D fingerprint matching.
In general, our work has three contributions:
@@
    The proposed method is the first monocular contactless fingerprint recognition algorithm that thoroughly operates in 3D space. The proposed method inputs a single 2D fingerprint image, and is able to output 3D features and 3D matching results. In contrast to previous contactless fingerprint methods [11][7][16][18] that only use 2D features during matching, our method maintains 3D features through the whole algorithm.
@@
    Our work makes use of the state-of-the-art graph neural network [21][22] that can fully represent the complicated relationships within structured data and has been successfully applied in fingerprint matching [23]. Our work further explores its application in contactless fingerprint and 3D fingerprint recognition.
@@
    We conduct thorough experiments on contactless fingerprint datasets with significant pose variations. Experiment results prove that our method can handle the pose variation and perspective distortion issues in contactless fingerprint recognition, and reaches the SOTA matching performances.
The following parts of this paper are: Section II is the literature review of recent contactless fingerprint algorithms; Section III introduces the proposed algorithm; Section IV is the experimental results on contactless fingerprint matching; and Section V is the conclusion.
II Related Work
There has been significant improvement in the area of contactless fingerprint recognition. Based on the types of feature extraction and matching approaches, the contactless fingerprint recognition algorithms can be generally categorized without/with 3D information.
II-A Contactless fingerprint recognition with only 2D feature
@@
Early studies [24][8][25][3] deal with contactless fingerprints the same as conventional contact-based fingerprints in feature extraction and matching methods. Further work [26][27][28] noticed the differences between contactless and contact-based fingerprints. The fingerprint quality and perspective distortion are the main issues of contactless fingerprints.
For contactless fingerprint recognition, specific feature extraction and matching methods are designed with the increasingly developed deep learning technique. Algorithms including segmentation [15][29], enhancement [19][15][16], minutiae extraction [30][11][7][31], and matching [14][19][16] are proposed to improve the recognition performance.
@@
The perspective distortion is another major issue in contactless fingerprint recognition. Early research [8][25][32][10] only used the central region or small rotation image of contactless fingerprints to decrease the impact of perspective distortion. A more general method is to train a distortion correction model [14][15][16][17] to stimulate the deformation grid, which is the same as in skin distortion rectification [13][33] in contact-based 2D fingerprint.
@@
In general, 2D-based contactless fingerprint recognition methods have made significant improvements, especially with the deep learning approach. But the 2D-based method fails to capture the 3D feature of contactless fingerprints, therefore has disadvantages in recognizing fingerprints of large 3D variations [11]. Also, learning the distortion field benefits the CL2C matching only, and is unsuitable for CL2CL matching.
II-B Contactless fingerprint recognition with 3D feature
@@
Recently, a few studies [11][20][18] have noticed the 3D information in contactless fingerprints and attempted to fuse 3D features into recognition. Usually, a finger-shaped model is estimated to stimulate the 3D finger. The reconstructed 3D shape is then used for pose correction [5][10][11] or 3D-2D unwarping [34][26][35][20][18]. However, the rotated or unwarped fingerprint is still a 2D image. The consequent feature extraction and matching algorithms are operated on the processed 2D contactless fingerprint. Therefore, the 3D information utilized here cannot be explicitly called ’3D feature’, as the fingerprint feature is yet 2D-based. The benefits of 3D information are not used in the final matching process.
@@
Several 3D fingerprint recognition algorithms have developed specific 3D features in 3D fingerprint matching, such as 3D minutiae [36][37], 3D ridge feature [38], and curvature [39][40]. However, the fingerprints used in these studies are 3D fingerprints acquired from specific 3D equipment like shape-from-shading [36][37] or shape-from-multiview [39], which are a bit different and more complicated than the contactless 2D fingerprints gathered by camera. Most existing image-based 3D fingerprint algorithms require two or more fingerprint captures and a calibrated camera system to restore the 3D information, which are too reluctant to be applied on the usual contactless fingerprint recognition scenes.
II-C Summary
In summary, using 3D features in contactless fingerprint recognition still needs further development. Current methods either lack 3D constraints in the recognition process, or do not fully utilize the 3D information during fingerprint feature and matching steps. Specialized 3D fingerprint methods are primarily redundant and inconvenient, and not fit for contactless fingerprints.
This paper proposes a monocular contactless fingerprint recognition with 3D feature and matching approaches. Our method can fully utilize the 3D information in contactless fingerprint recognition with an efficient and robust processing approach.
III Methods
@@
Fig. 1 shows the framework of the proposed method. Our method mainly consists of three steps: 1) the preprocessing step that segments, enhances, and unifies the scale and pose angles of fingerprint images [15][18]; 2) the feature extraction step that goes through a multi-task network to output the 3D shape and minutiae features; 3) the matching step that uses the 3D minutiae map and a 3D graph matching network [23] to output the matching scores. Although the input contactless fingerprint is 2D, our method can recover the 3D shape model, and therefore conduct a 3D feature matching algorithm.
III-A Preprocessing
@@
Due to the complicated situations of contactless fingerprints, it is essential to minimize the variations within training data to strengthen the robustness of the proposed network. Similar to previous studies [15][16][18], we primarily standardize contactless fingerprint images from three perspectives: contrast, pose, and ridge frequency.
Firstly, as contactless fingerprint images are often captured under different lighting conditions, it is crucial to unify the overall brightness range of the images. For datasets like UWA[41] and CFPose[11], we initially selected region with grayscale values greater than 15 as the foreground region. The maximum convex hull of these regions is identified as a mask, followed by the application of histogram equalization to normalize the contrast of the segmented images while minimizing the effect of noise.
Secondly, the finger’s shape is also captured during contactless fingerprint capture. The finger’s shape allows us to directly obtain the centerline of the finger through the contour in the preprocessing stage. Rotating the finger’s centerline to a vertical position eliminates rotations along the $z$-axis, providing a certain level of normalization.
Lastly, the scale of contactless fingerprint capture varies significantly, leading to substantial variations in the frequency of captured fingerprint ridge lines. It is necessary to normalize ridge frequency, which could affect fingerprint recognition. Thus, the ridge frequency in the central region of the fingerprint is balanced to 10 pixels across fingerprint images, avoiding affecting the subsequential feature extraction and matching steps.
III-B Feature Extraction
The feature extraction network takes a preprocessed contactless fingerprint as the input, and then outputs the estimated gradient, minutiae, ridge period, and orientation map. Fig. 2 shows examples of two sets of fingerprint features, including orientation, period, minutiae, and gradient. Since the gradient is a two-dimensional vector, it is represented as the combination of scalar X and Y.
@@ III-B Feature Extraction
It should be noted that the direct results from the network are both in 2D. The estimated surface gradient is used for 3D reconstruction [18] to get a 3D finger shape. Then, the minutiae and orientation are recomputed according to the reconstructed 3D shape, and we can finally get the 3D minutiae result.
III-B1 Network Structure
Figure 2: Examples of the fingerprint features.
Figure 3: Pipeline of the proposed feature extraction network.
@@ III-B1 Network Structure
Fig. 3 shows the pipeline of the proposed feature extraction network. The network inputs the preprocessed fingerprint and segmentation mask from the above preprocessing step and outputs the estimated gradient, minutiae, ridge period, and orientation. This network is an adaptation from [18] with an extra minutiae branch, which is inspired by FingerNet [42]. This network has multi-task branches, which has been proven in [18] that a multi-task network performs better than a single task. In [18], the estimated orientation and ridge period serve only as a middle layer output and do not contribute to the final matching process. In our method, the output minutiae, orientation, and 3D shape are all combined to serve as a set of 3D features for the following matching step.
Figure 4: Details of the proposed feature extraction network structure.
The detail of the feature extraction network is shown in Fig. 4. The left branch is the ridge feature extraction branch, and the right is the minutiae branch. The ridge feature branch is the same as in [18], where network details can be found. The right minutiae branch has a feature extraction part of three convolution and max-pooling layers, and the following steps corresponding to four minutiae outputs as in [42].
The minutiae map contains four outputs [42]: minutiae score map, precise X and Y location map, and minutiae orientation map. The minutiae score map is filtered by a threshold to get the minutiae point, then it adds the precise X and Y location map to get the final minutiae location. The minutiae orientation is the same as the left orientation branch, only expanding $180^{\circ}$ to $360^{\circ}$.
III-B2 Loss Definition
@@
The orientation loss $\mathcal{L}_{\text{Ori}}$, period loss $\mathcal{L}_{\text{Ped}}$, and gradient loss $\mathcal{L}_{\text{Grad}}$ are the same as in [18], where detailed descriptions can be found. Here, we simply further explain the minutiae loss $\mathcal{L}_{\text{Minu}}$.
$\mathcal{L}_{\text{Ori}}=\text{CE}(o(x,y),o^{*}(x,y))+\alpha\text{Coh}(o(x,y))$ (1)
$\mathcal{L}_{\text{Ped}}$ $=-\frac{1}{|M|}\sum_{M}(p(x,y)-p^{*}(x,y))^{2}+\beta\|\nabla p(x,y)\|^{2}$ (2)
$\mathcal{L}_{\text{Grad}}$ $=-\frac{1}{\sum_{M}{w(x,y)}}\sum_{M}w(x,y)\|g(x,y)-g^{*}(x,y)\|^{2}$ (3)
$+\gamma\|\nabla g(x,y)\|^{2}$
As the minutiae branch has four outputs, the minutiae loss $\mathcal{L}_{\text{Minu}}$ is the sum of four parts corresponding to minutiae score loss $\mathcal{L}_{\text{M}1}$, X location loss $\mathcal{L}_{\text{M}2}$, Y location loss $\mathcal{L}_{\text{M}3}$, and minutiae orientation loss $\mathcal{L}_{\text{M}4}$. As described in [42], both $\mathcal{L}_{\text{M}1}$, $\mathcal{L}_{\text{M}2}$, and $\mathcal{L}_{\text{M}3}$ use the 0-1 heatmap as the ground truth, where the correct minutiae location is set to 1, and other points are set to 0. The losses would be the cross entropy between the estimated heatmap and ground-truth heatmap. As for minutiae orientation loss $\mathcal{L}_{\text{M}4}$, it is also a cross entropy between the output orientation and the ground truth, which is the same as in orientation loss $\mathcal{L}_{\text{Ori}}$. The final minutiae loss would be:
$\mathcal{L}_{\text{Minu}}=\mu_{1}\mathcal{L}_{\text{M}1}+\mu_{2}\mathcal{L}_{\text{M}2}+\mu_{3}\mathcal{L}_{\text{M}3}+\mu_{4}\mathcal{L}_{\text{M}4}.$ (4)
$\mu_{1}$-$\mu_{4}$ are the weights of these minutiae losses.
@@
$\lambda_{1}$-$\lambda_{4}$ are the weights of these four losses. All parameters used in the loss function are listed in Table I.
TABLE I: Parameters used in the loss function of feature network
Parameter $\alpha$ $\beta$ $\gamma$ $\sigma$
Value 1 1 1 0.5
Parameter $\mu_{1}$ $\mu_{2}$ $\mu_{3}$ $\mu_{4}$
Value 300 15 15 15
Parameter $\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$
Value 1 60 300 1
III-B3 Training Data
We use the UWA contactless fingerprint database [41] to generate the ground-truth training data of the feature extraction network. This database contains contactless fingerprints from different viewing angles, which is beneficial for constructing a robust recognition method for various fingerprint poses. Also, this database was acquired by the Surround Imager [9], which uses three calibrated cameras to reconstruct a 3D finger shape using the shape-from-silhouette algorithm [43]. The detailed 3D finger reconstruction algorithm can be found in the supplementary material of [18].
For fingerprints in this database, we first preprocess the contactless fingerprint image using the method described in Section III-A. Then, the orientation field, period map, and minutiae map are extracted by the commercial software VeriFinger 13.0 [44]. The 3D shape is computed using the method in [18], and the gradient of the finger shape can be obtained using numerical differential. The generated gradient, orientation, period, and minutiae are used as ground-truth data for training the network.
UWA database has 8,958 contactless fingerprints, and each finger has 3 viewing angles $\times$ 2 images. Following [15], we use the last 2,988 fingerprints as training data, and the top 5,970 for matching experiments. For the 2,988 fingerprints, we crop 5,506 patches of size $512\times 512$ to train, validate, and test the network. The sizes of all outputs (gradient, orientation, period, minutiae map) are $1/8$ of the input size.
III-C 3D Graph Matching
(a) 3D minutiae
(b) 3D minutiae on 3D fingers
Figure 5: Examples of 3D fingerprints and 3D minutiae.

### v2

# Improving Contactless Fingerprint Recognition with Robust 3D Feature Extraction and Graph Embedding

@@ Abstract
Contactless fingerprint has gained lots of attention in recent fingerprint studies. However, most existing contactless fingerprint algorithms treat contactless fingerprints as 2D plain fingerprints, and still utilize traditional contact-based 2D fingerprints recognition methods. This recognition approach lacks consideration of the modality difference between contactless and contact fingerprints, especially the intrinsic 3D features in contactless fingerprints. This paper proposes a novel contactless fingerprint recognition algorithm that captures the revealed 3D feature of contactless fingerprints rather than the plain 2D feature. The proposed method first recovers 3D features from the input contactless fingerprint, including the 3D shape model and 3D fingerprint feature (minutiae, orientation, etc.). Then, a novel 3D graph matching method is proposed according to the extracted 3D feature. Additionally, the proposed method is able to perform robust 3D feature extractions on various contactless fingerprints across multiple finger poses. The results of the experiments on contactless fingerprint databases show that the proposed method successfully improves the matching accuracy of contactless fingerprints. Exceptionally, our method performs stably across multiple poses of contactless fingerprints due to 3D embeddings, which is a great advantage compared to 2D-based previous contactless fingerprint recognition algorithms.
1 Introduction
@@
Compared with traditional contact-based fingerprint, which relies on pressing finger skin against sensors to gather fingerprint images, contactless fingerprint provides a touchless approach to acquiring fingerprint images. Since the collection process is touch-free, the gathered fingerprints are of no skin distortion, which is a crucial problem in contact-based fingerprint matching [bazen2003matching]. The collection process is more hygienic and convenient and can even be done with mobile phone cameras [stein2012fingerphoto][sankaran2015smartphone].
@@
Despite these advantages, contactless fingerprint recognition still meets challenges on modality differences with contact-based fingerprints [tan2021minu]. The contactless fingerprint image is acquired by a camera-like imaging system [lee2006study][parziale2006surround] in 3D space, which differs from traditional contact-based 2D sensors. The finger surface curvature is incorporated into the imaging result, which leads to evident perspective distortion on contactless fingerprints [labati2013contactless][tan2020towards]. The distortion causes changes in orientation, period, and minutiae location, thus degrading recognition performances [jain1997line].
Figure 1: Compared with previous methods that only uses 2D fingerprint feature, our method extracts 3D features of contactless fingerprints and encodes them via a 3D graph neural network for fingerprint matching.
@@
To overcome perspective distortion and accurately match contactless fingerprints, researchers have proposed methods to deal with this issue. A significant approach is to treat perspective distortion similar to elastic skin distortion [si2015detection] in contact-based 2D fingerprints. A transformation network [lin2018matching][Dabouei2019][grosz2022contact][dong2023syn] is trained to fit the distortion, thus rectifying the fingerprint to the undistorted level. This approach is generally used in contactless-to-contact (CL2C) matching, since it treats a contactless fingerprint as a distorted 2D fingerprint, similar to a contact-based 2D fingerprint.
@@
As for contactless-to-contactless (CL2CL) matching, several studies have further considered the intrinsic 3D information of contactless fingerprints. A finger shape model [tan2020towards][tan2021minu][cui2023monocular] is recovered from the original contactless fingerprint. Then, all the contactless fingerprints are rotated to the same pose (usually the front pose) in 3D space to eliminate pose variations for improving the matching accuracy.
@@
Although significant progress has been made in contactless fingerprint recognition, the modality characteristics of contactless fingerprints have yet to be fully studied. Most algorithms [lin2018matching][Dabouei2019][lin2019cnn][grosz2022contact][dong2023syn] treat contactless fingerprints as plain 2D fingerprint images, therefore neglecting the 3D information hidden in contactless fingerprints. Some studies further consider 3D information. However, the recovered 3D shape is only used for pose correction [labati2013contactless][tan2020towards] or unwarping [Sollinger2021][cui2023monocular]. The rotated or unwarped contactless fingerprint is still treated as a 2D image for the following feature extraction and matching steps.
@@
In this paper, we propose a novel contactless fingerprint recognition method that uses 3D information in contactless fingerprints. As shown in Fig. 1, unlike previous methods that only extracted 2D features from contactless fingerprint images, the proposed method takes a contactless fingerprint as input, and then extracts its 3D feature, including the shape model, 3D minutiae feature, etc. Based on those features, we can reconstruct the feature points in 3D space and extract 3D feature embeddings. These 3D features are finally used for 3D fingerprint graph matching.
In general, our work has four contributions:
@@
    This proposed method innovatively extracts features for contactless fingerprint recognition within a 3D space. The proposed method inputs a single 2D fingerprint image, and is able to output 3D features and 3D matching results. In contrast to previous contactless fingerprint methods [tan2020towards][tan2021minu][grosz2022contact][cui2023monocular][dong2023syn][shi2022towards] that only use 2D features during matching, our method maintains 3D features throughout the algorithm.
@@
    Our method improves the monocular depth estimation network [cui2023monocular], enhances the depth prediction accuracy of previous methods, and addresses the issue of poor generalization of previous methods to various contactless fingerprint datasets.
@@
    Our work makes use of the commonly used graph neural network [li2019deepgcns][Wu2021Survey] that can represent the complicated relationships within structured data and has been successfully applied in fingerprint matching [su2023gnn]. Our work further explores its application in contactless fingerprint and 3D fingerprint recognition.
  • 4)
@@
    We conduct thorough experiments on contactless fingerprint datasets with significant pose variations. Experiment results on several datasets prove that our method can handle the pose variation and perspective distortion issues in contactless fingerprint recognition, and reaches the SOTA matching performances.
Figure 2: The pipeline of the proposed method. Our method first extracts 3D features (3D shape, 3D minutiae feature, deep features, etc.) from the input contactless 2D fingerprint, then uses the extracted 3D features to conduct matching both in 2D space with unwarping and in 3D space by extracting fixed length graph embedding.
2 Related Work
2.1 Contactless fingerprint recognition with only 2D feature
@@
Early studies [song2004new][lee2006study][hiew2007touch][parziale2008touchless] dealt with contactless fingerprints the same as conventional contact-based fingerprints in feature extraction and matching methods. Further work [chen2006touchless][zhou2013performance][zhou2014performance] noticed the the fingerprint quality and perspective distortion issues of contactless fingerprints and made several improvements.
For contactless fingerprint recognition, specific feature extraction and matching methods are designed with the deep learning technique. Algorithms including segmentation [Dabouei2019][Malhotra2020], enhancement [lin2019cnn][Dabouei2019][grosz2022contact], minutiae extraction [lin2017multi][tan2020towards][tan2021minu][zhang2021multi], and matching [lin2018matching][lin2019cnn][grosz2022contact] are proposed to improve the recognition performance. Shi et al. [shi2022towards] adopted graph neural networks for contactless fingerprint matching, just as we did. However, they did not construct a graph in 3D space. Moreover, their method ultimately matched the minutiae of contactless fingerprints, whereas our method encodes three-dimensional fingerprints into fixed-length features.
@@
The perspective distortion is another major issue in contactless fingerprint recognition. Early research [lee2006study][hiew2007touch][kumar2011contactless][labati2013contactless] only used the central region or small rotation image of contactless fingerprints to decrease the impact of perspective distortion. A more general method is to train a distortion correction model [lin2018matching][Dabouei2019][grosz2022contact][dong2023syn] to stimulate the deformation grid, which is the same as in skin distortion rectification [si2015detection][Dabouei2018] in contact-based 2D fingerprint.
@@
In general, 2D-based contactless fingerprint recognition methods have made significant improvements, especially with the deep learning approach. But the 2D-based method fails to capture the 3D feature of contactless fingerprints, therefore has disadvantages in recognizing fingerprints of large 3D variations [tan2020towards]. Also, learning the distortion field benefits the CL2C matching only, and is unsuitable for CL2CL matching.
2.2 Contactless fingerprint recognition with 3D feature
@@
Recently, a few studies [tan2020towards][Sollinger2021][cui2023monocular] have noticed the 3D information in contactless fingerprints and attempted to fuse 3D features into recognition. Usually, a finger-shaped model is estimated to stimulate the 3D finger. The reconstructed 3D shape is then used for pose correction [stein2012fingerphoto][labati2013contactless][tan2020towards] or 3D-2D unwarping [fatehpuria2006acquiring][chen2006touchless][Shafaei2009new][Sollinger2021][cui2023monocular]. However, the rotated or unwarped fingerprint is still a 2D image. The consequent feature extraction and matching algorithms are operated on the processed 2D contactless fingerprint. Therefore, the 3D information utilized here cannot be explicitly called ’3D feature’, as the fingerprint feature is yet 2D-based. The benefits of 3D information are not used in the final matching process.
@@
Several 3D fingerprint recognition algorithms have developed specific 3D features in 3D fingerprint matching, such as 3D minutiae [kumar2013towards][lin2017tetrahedron], 3D ridge feature [yin2021fingerprint], and curvature [liu2015study][galbally2017full]. However, the fingerprints used in these studies are 3D fingerprints acquired from specific 3D equipment like shape-from-shading [kumar2013towards][lin2017tetrahedron] or shape-from-multiview [liu2015study], which are a bit different and more complicated than the contactless 2D fingerprints gathered by camera. Most existing image-based 3D fingerprint algorithms require two or more fingerprint captures and a calibrated camera system to restore the 3D information, which are too reluctant to be applied on the usual contactless fingerprint recognition scenes.
3 Methods
@@
Fig. 2 shows the framework of the proposed method. Our method consists of three main steps: 1) the preprocessing step that segments, enhances, and unifies the scale and pose angles of fingerprint images [Dabouei2019][cui2023monocular]; 2) the feature extraction step that goes through a multitask network to output the 3D shape and minutiae features; 3) the matching step that uses the 3D feature and a 3D graph matching network [su2023gnn] to output the matching scores. Although the input contactless fingerprint is 2D, our method can recover the 3D shape model, and therefore conduct a 3D feature matching algorithm.
The score of 3D graph matching offers a new perspective for 2D fingerprint matching and can effectively enhance the performance of contactless fingerprint recognition. Meanwhile, we use our newly proposed monocular depth estimation network with better generalization to unwarp contactless fingerprints, correcting perspective distortion and matching them in a 2D space. By combining the two matching approaches, our method can effectively boost the performance of contactless fingerprint recognition.
3.1 Preprocessing
@@
Due to the complicated situations of contactless fingerprints, it is essential to minimize the variations within training data to strengthen the robustness of the proposed network. Similar to previous studies [Dabouei2019][grosz2022contact][cui2023monocular][dong2023syn], we primarily standardize contactless fingerprint images from four perspectives: segmentation, contrast, pose and ridge frequency.
In fingerprint recognition tasks, attention is usually focused only on the first phalanx. Similar to Dong and Kumar [dong2023syn], we employ SOLOv2 [wang2020solov2] for the segmentation of the first phalanx. We annotated the first 912 images in the ZJU dataset for SOLOv2 training and validation, and segmented the remaining part of the ZJU dataset, as well as the UWA [zhou2014benchmark] and CFPose [tan2020towards] datasets. We then employ the same method as in [cui2023monocular] to adjust fingerprint contrast, orient the fingerprints correctly, and unify ridge frequency.
Figure 3: The detailed pipeline of the proposed 3D feature network, which extracts 3D features from the input contactless 2D fingerprint. The network outputs a 128-dimensional deep feature along with the depth and minutiae feature for the following graph embedding. Each decoder (period, orientation, gradient, minutiae) contains 2 inputs and 1 output: (a) The deepest features of Lite-Mono Encoder as input. (b) The middle-level features of Lite - Mono Encoder that are used as input. (d) Final output. And (c) is the intermediate output only in period and orientation decoder.
3.2 Fingerprint Feature Extraction
@@
The 3D feature network takes a preprocessed contactless fingerprint as input, and then outputs the surface gradient, minutiae location, ridge period and orientation. It should be noted that the direct results from the network are in 2D. The estimated surface gradient is used for 3D reconstruction [cui2023monocular] to get a 3D finger shape. Then, the minutiae and orientation are recomputed according to the reconstructed 3D shape, and we can finally get the 3D minutiae result.
@@
Fig. 3 shows the pipeline of the proposed feature extraction network. The network inputs the preprocessed fingerprint and segmentation mask from the above preprocessing step, and outputs the estimated gradient, minutiae, ridge period, and orientation. This network is an adaptation from [cui2023monocular] with an extra minutiae branch, which is inspired by FingerNet [fingernet2017]. This network has multitask branches, which has been proven in [cui2023monocular] that a multitask network performs better than a single task.
However, the method in [cui2023monocular] still cannot perform depth estimation well on various contactless fingerprint datasets under different acquisition conditions, and the resulting depths are often not smooth enough. Therefore, we improved this network by adopting the Lite-Mono encoder from [zhang2023lite] and fine-tuning it with pre-trained parameters, which enhanced the generalization ability of the previous network. Compared to standard Litemono encoder, our Litemono encoder only outputs feature maps of two scales, with 80 and 128 channels respectively. For each decoder, the 128-channel feature map serves as the input to the first convolutional layer, and the 80-channel feature map is concatenated with the output of the second convolutional layer to form the input to the third convolutional layer.
The four decoders (period, orientation, gradient, minutiae) process different output features, and are of the same structure with the only difference being the number of output channels. The structure of decoders is shown in the middle of the right side of Fig. 3, which is a series of convolution and upsampling layers. The input arrow to the second downsampling layer represents the skip connection of the Lite-Mono layer. The first output arrow represents the feature of the previous layer, and the feature of the direction field and period map is used for predicting the gradient. The second output arrow represents the final prediction result.
The minutiae branch uses the Local MinNet from the minutiae extraction network in [feng2023detecting] for minutiae localization, as Lite-Mono Encoder [zhang2023lite] is already capable of extracting global information from the input fingerprint effectively. Finally, the output minutiae are obtained by multiplying the local prediction results from the minutiae branch with the global prediction results from the depth estimation network. The minutiae map branch only outputs the minutiae location. The minutiae orientation is the same from the orientation branch, only expanding $180^{\circ}$ to $360^{\circ}$.
@@
The orientation loss $\mathcal{L}_{\text{Ori}}$, period loss $\mathcal{L}_{\text{Ped}}$, and gradient loss $\mathcal{L}_{\text{Grad}}$ are the same as in [cui2023monocular], where detailed descriptions can be found. We construct the minutiae location map in the same manner as [feng2023detecting] and adopt their proposed Dynamic End-to-End Loss Design as our $\mathcal{L}_{\text{Minu}}$.
@@
$\lambda_{1}$-$\lambda_{4}$ are the weights of these four losses. All parameters used in the loss function are same as Cui et al. [cui2023monocular] except $\lambda_{4}=1000$.
We upsample the output gradients, minutiae location maps, orientations, and 128-dimensional depth feature maps to the original image size. The gradients are used to reconstruct depth, while orientations and minutiae maps are transformed into 2D minutiae using the method from [feng2023detecting]. Finally, gradients, depth, 2D minutiae, and depth feature maps are combined to construct the 3D fingerprint features for subsequent graph embedding.
3.3 3D Feature Graph Embedding
3D Minutiae Feature. The direct outputs of the feature extraction network are 2D. They need to be transformed into 3D space to fit for the 3D graph matching. Compared with 2D minutiae, 3D minutiae have an extra dimension, which needs to be calculated out.
The 3D depth $z$ is recovered from the gradient results $g_{x}$ and $g_{y}$ using numerical integration in [cui2023monocular], as the gradient is the differential of the surface depth. The calculated finger depth represents the $z$ coordinate, and is combined with the original image pixel $(x,y)$ to get a 3D coordinate $\mathbf{p}=(x,y,z)$. The direction of 3D minutiae can be obtained through a two-step rotation in the local coordinate system. The minutiae direction $\theta$ is the angle of the 3D minutiae in the x-y plane, which is the angle of the first step rotation. The gradients $g_{x}$ and $g_{y}$ represent the changes in axis $z$ respectively when $x$ and $y$ changes by one pixel. Therefore, the change $d_{\theta}$ in one pixel along the $\theta$ direction is
$d_{\theta}=\cos\theta\frac{g_{x}}{\sqrt{g_{x}^{2}+1}}+\sin\theta\frac{g_{y}}{\sqrt{g_{y}^{2}+1}}$ (2)
This change can represent the tangent value of the angle $\varphi$ in the second step, that is, $\varphi=\arctan d_{\theta}$. Thus, based on the spherical coordinate system, the final direction of the 3D minutia is:
$\begin{bmatrix}d_{x}\\ d_{y}\\ d_{z}\end{bmatrix}$= $\begin{bmatrix}cos\theta&-\sin\theta&0\\ sin\theta&\cos\theta&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}cos\varphi&0&-\sin\varphi\\ 0&1&0\\ sin\varphi&0&\cos\varphi\end{bmatrix}$ (3)
$\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}1\\ 0\\ 0\end{bmatrix}=\begin{bmatrix}\cos\theta\cos\varphi\\ \sin\theta\cos\varphi\\ \sin\varphi\end{bmatrix}$
It is important to note that the Euler angle transformation used here is based on a local coordinate system, and the order of matrix multiplication is the opposite of that in a global coordinate system.